If an angle of a triangle remains unchanged but each of its two including sides is doubled, then the new area divided by the old area is
(a) 2 (b) 3 (c) 4 (d) 5 (e) more than 6

Answer by Theo(2967)

Answer is (c) = 4.

click HERE for a picture of the problem.

Basically, when you double the two sides with the angle in between them, then you double the altitude, and when you double the altitude, you quadruple the area because.

A = 1/2 * b * h where:

b is the base = one of the included sides, and:

h is the altitude.

b is doubled and h is doubled to get:

A = 1/2 * 2 * b * 2 * h = 1/2 * 4 * b * h